OpenStudy (anonymous):

Todd currently has $5,000. What was the value of his money four years ago if he's earned 4 percent interest each year?

OpenStudy (anonymous):

future value=p x(1+i)t present value=p/(1+i)t

OpenStudy (cwrw238):

i'm assuming its simple interest right?

OpenStudy (anonymous):

yep that's right

OpenStudy (cwrw238):

if his original sum was p dollars the p + 4*(0.04p) = 5000 1 .16p = 5000 p = 5000/1.16 = 4310.34 dollars

OpenStudy (anonymous):

nice, now this completely makes sense.... Thanks so much I have another one if you cold plz help If you invest $2,000 today in a bank that gives you a 4 percent annual interest rate, which of these items can you buy in seven years?

OpenStudy (anonymous):

a home theater system worth 3000 a designer dress worth 2700 or a mountain bike worth 2500 ?

OpenStudy (cwrw238):

work out how much money you will get in 7 years 2000 +interest for 7 years = 2000 + 7*0.04* 2000

OpenStudy (cwrw238):

what do you make that?

OpenStudy (anonymous):

I got 2,560

OpenStudy (cwrw238):

right - so im afraid you cant buy your designer dress...

OpenStudy (anonymous):

aww, drat,, j/k so I guess I'll at least be able to buy the mountain bike

OpenStudy (cwrw238):

but you can get fit with a mountain bike

OpenStudy (anonymous):

true

OpenStudy (cwrw238):

lol

OpenStudy (anonymous):

you are funny

OpenStudy (anonymous):

If you save $3,000 at an interest rate of 14 percent per year, how much will you have at the end of six years?

OpenStudy (anonymous):

3000 plus 6 times 0.14 x 3,000 right?

OpenStudy (anonymous):

I need help I don't think this is right

OpenStudy (anonymous):

A(t)=p(1+rt) use this formula